OpenStudy (anonymous):

Help me please!!!... I am trying to match this equation with its graph x^3/x^2-1 but I can't figure out the x and y intercept in this equation.

OpenStudy (zzr0ck3r):

\(y=\frac{x^3}{x^2-1}\)?

OpenStudy (zzr0ck3r):

what is y when x=0? what is x when y = 0? hint: they are the same number

OpenStudy (anonymous):

f(x)=x^3/x^2-1

OpenStudy (zzr0ck3r):

ok all you did was repeat it.....is it the one i put?

OpenStudy (zzr0ck3r):

yes or no:)

OpenStudy (anonymous):

Ok

OpenStudy (zzr0ck3r):

what?

OpenStudy (zzr0ck3r):

lol

OpenStudy (anonymous):

I don't get what you mean @zzr0ck3r

OpenStudy (zzr0ck3r):

it is hard to tell what equation you are asking about......I ask you a yes or no question, to which you reply "ok". please read what I wrote and answer YES or NO.

OpenStudy (anonymous):

yes

OpenStudy (zzr0ck3r):

ok what is y when x = 0?

OpenStudy (anonymous):

0?

OpenStudy (zzr0ck3r):

so the y intercept is (0,0)

OpenStudy (zzr0ck3r):

what is x when y = 0?

OpenStudy (zzr0ck3r):

a fraction can only = 0 if the top part equals 0 so what is the answer to x^3=0?

OpenStudy (anonymous):

0

OpenStudy (zzr0ck3r):

sso the x intercept is (0,0)